OpenStudy (anonymous):

Choose the correct simplification of the expression (8x2)3.

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \left(8x^2 \right )^3 = 8^3 \left(x^2 \right )^3\]

OpenStudy (unklerhaukus):

\[\left(ab^m\right)^n=a^nb^{mn}\]

OpenStudy (richyw):

\(8^3x^6\)